OpenStudy (anonymous):

In an examination it is required to get 40% of the aggregate marks to pass. A student gets 261 marks and is declared failed by 4% marks. What are the maximum aggregate makrs a student can get?

OpenStudy (kropot72):

261 marks is (40 - 4) = 36% of the aggregate marks. Maximum aggregate marks = 261/0.36 Now can you solve it?

OpenStudy (anonymous):

36 % = 261 => 100 % = 261 * 100/36= ?

OpenStudy (kropot72):

jaikiran323 can you use a calculator?